# A Deep Learning Framework for Disentangling Triangle Singularity and Pole-Based Enhancements

### v2

@@ III Uniformization of $S$-matrix
Enhancements observed in scattering cross-sections most often correspond to poles in the scattering matrix. Different $S$-matrix parametrizations can be used to analyze a given experimental data. Perhaps, the most popular one is the $K$-matrix parametrization since the unitarity of the $S$-matrix can be automatically satisfied. However, identifying the location and Riemann sheet of the poles requires solving a polynomial in channel momenta. As such, their locations cannot be controlled to ensure a well-distributed representation of different possible pole locations. Moreover, it is usually the case that a pole can be placed on the desired Riemann sheet in the $K$-matrix approach. However, these controlled poles are usually accompanied by shadow poles which may appear in the physical Riemann sheet. Such pole configuration will result in a violation of causality and must not be included in our training dataset. It is, therefore, more practical to use a parametrization in which poles can be controlled directly. It turns out that this problem can be addressed by a more rigorous approach called uniformization Santos and Sombillo (2023).
@@
where $\epsilon=m_{1}+m_{2}$ is the threshold, and $\mu$ is the reduced mass. Notice that even though we show $p$ in terms of $\sqrt{s}$, there are still branch point singularities due to the existence of square root cross terms. To resolve this, we introduce the uniformized momentum $q_{i}$ to simplify the scattering energy, giving us
$s=q^{2}+\epsilon^{2}.$ (11)
For two-channel scattering, we have
$s=q_{1}^{2}+\epsilon_{1}^{2}=q_{2}^{2}+\epsilon_{2}^{2}.$ (12)
The two-channel scattering can be described by a $2\times 2$ uniformized $S$-matrix whose elements are constructed using a Jost function $D$ and are given by
$S_{11}(q_{1},q_{2})$ $=\dfrac{D(-q_{1},q_{2})}{D(q_{1},q_{2})};$ (13)
@@
which can be derived by utilizing the two expressions in equation (12). Hence, we arrive at a Jost function with the form
@@
The factor $(\omega-\omega_{\text{pole}})$ produces the poles in the $S$-matrix. $1/\omega^{2}$ prohibits very energetic initial states from having no final states after the interaction as this is unphysical. The pole $\omega_{\text{reg}}$ ensures that the diagonal elements go to unity ($S_{ii}\rightarrow 1$) when the initial states have very large momentum ($\omega\rightarrow\infty$), which is a well-known property in potential scattering theory Taylor (1972). It is controlled such that it is far from the relevant pole and does not affect the interpretation of the enhancements. To satisfy hermiticity below the lowest threshold, there must be a pole in the negative complex conjugate of $\omega_{\text{pole}}$, that is why we have the factors ($\omega+\omega_{\text{pole}}^{*}$)($\omega+\omega_{\text{reg}}^{*}$).
As such, by using uniformization, one has the freedom to adjust the position of the pole and its regulator to their desired Riemann sheet while preserving causality. In two-channel scattering, there are four Riemann sheets of the complex energy plane. The first Riemann sheet named [$tt$]-sheet is called the physical sheet where the scattering region is located. There should be no poles in the physical sheet to obey causality van Kampen (1953a, b). The other three Riemann sheets, namely [$bt$], [$bb$], and [$tb$]-sheets, are the unphysical sheets where poles located can be associated with the presence of an unstable quantum state in the reaction involved. One can say that the pole is located in the [$bt$]-sheet when the peak is below the second threshold, and a pole is located in the [$tb$]-sheet when the peak is located exactly at the threshold. These pole structures, and their manifestation on the physical region shown in Figure 4, are useful in the description of the hadron’s internal structure as demonstrated in the pole-counting argument by Morgan and Pennington Morgan (1992); Morgan and Pennington (1991, 1993). The location of the Riemann sheets is shown in Table 1. The pole regulator has the form $\omega_{\text{reg}}=\exp(-i\pi/2)/|\omega_{\text{pole}}|$ where the phase factor forces the regulator to be in either the [$bb$]-sheet or the [$tb$]-sheet below the lowest threshold.
Figure 4: Expected line shapes arising from isolated poles in one of the unphysical Riemann sheets. The yellow dotted line is the second threshold. The energy poles were generated with real part exactly at the second threshold and imaginary part of $-10$ MeV.
@@
The $J/\psi p$ scattering is a coupled-channel scattering problem where the first channel is set to always be $J/\psi p$ and the second channel is $\Sigma_{c}^{+}\bar{D}^{0}$ as this is the prominent threshold near the $P_{\psi}^{N}(4312)^{+}$ signal. One can construct its uniformized $S$-matrix where the elements $S_{11}$ describes the transition of $J/\psi p\rightarrow J/\psi p$, $S_{12}$ for $J/\psi p\rightarrow\Sigma_{c}^{+}\bar{D}^{0}$, $S_{21}$ for $\Sigma_{c}^{+}\bar{D}^{0}\rightarrow J/\psi p$, and $S_{22}$ for $\Sigma_{c}^{+}\bar{D}^{0}\rightarrow\Sigma_{c}^{+}\bar{D}^{0}$. A compact state arises when there are poles both in the [$bt$]-sheet and [$bb$]-sheet, corresponding to the mechanism in Figure 3(c). By placing two identical energy poles in the [$bt$]-sheet and [$bb$]-sheet (i.e. the poles are equidistant from the relevant threshold), one can interpret the fitted enhancement as a pure compact state same as fitting a Breit-Wigner amplitude Baru et al. (2004). On the other hand, a quasi-bound state interpretation is possible whenever there is an isolated pole in the $[bt]$-sheet with strong coupling to the higher channel, corresponding to Figure 3(d). However, this interpretation should be examined carefully since a virtual state of the higher channel can move to the second Riemann sheet due to strong channel-coupling Fernández-Ramírez et al. (2019); Frazer and Hendry (1964); Pearce and Gibson (1989). It must be pointed out that bound and virtual states fall under the same general classification of hadronic molecules as demonstrated in Matuschek et al. (2021). In any case, any interpretation should be supported by the corresponding pole trajectory, which can only be done in a model-dependent way Badalyan et al. (1982); Baru et al. (2004); Hanhart et al. (2014).
@@
Nonetheless, through uniformization, one can easily construct pole-based amplitudes that correspond to the interpretations discussed. This is demonstrated in Figure 5. We assign an isolated pole with energy (in units of MeV) at $4317.49-7j$ in the [$bt$]-sheet to construct an amplitude representing a possible hadronic molecule, while we assign one pole each at $4311.90-7j$ in both the [$bt$] and [$bb$]-sheets representing a compact pentaquark state. An amplitude is generally complex-valued with a real and imaginary part, but for our purposes, it is sufficient to consider its modulus square which corresponds to the intensity. Therefore, plotting the $|T_{11}|^{2}$ ($T$-matrix) contribution using the uniformized poles and following equation (8), we see that the two similar line shapes fit the LHCb experimental data well, both with a prominent peak at 4312 MeV. If we compare Figure 5 with Figure 2, we notice that the pole-based enhancements have a seemingly narrower width (FWHM) compared to that produced by a triangle singularity, and can fit the experimental error bars better especially near the $\Sigma_{c}^{+}\bar{D}^{0}$ threshold. We take note of our observations as we move on to constructing the training dataset for our DNN.
@@ IV.1 Generation of training datasets
Our goal is to construct a DNN model that accepts two-dimensional line shapes as input features and maps them to output targets corresponding to the possible classifications. Since we will eventually apply the trained DNN on interpreting the experimental data, the training dataset must follow the general characteristics of the experimental data (similar to our demonstrations in Figures 2 and 5). Thus, it is important to take into consideration the limited energy resolution of the data arising from the statistics of the particle detector. In the LHCb data LHCb Collaboration (2019), the energy axis is divided into bins of 2 MeV each. For our training dataset, we can generate amplitudes on randomly spaced energy points, where one representative energy is randomly chosen for each bin. We use a uniform distribution for this task. Moreover, we restrict our region of interest to the energy region of 4200 MeV to 4350 MeV, in order to avoid other prominent structures aside from the $P_{\psi}^{N}(4312)^{+}$. This region consists of 75 energy bins, implying that our line shapes shall have 75 points each for the horizontal and vertical components. This totals to 150 input nodes, as presented in Figure 6.
@@
For the output targets, we create labels for each of the possible classifications, as listed in Table 3. The triangle singularity takes up the first label, while the rest are possible pole configurations. As previously stated, 1 pole in [$bt$] may be associated to the hadronic molecule interpretation, 1 pole in [$tb$] corresponds to a virtual state, and 1 pole in [$bt$] and 1 pole in [$bb$] can accommodate a compact pentaquark interpretation. Refer to section III for the detailed discussion. It is possible to include more sophisticated pole configurations, such as considering 2 or more poles in each of the sheets as done in Sombillo et al. (2021). However, for the purposes of our task which is mainly to distinguish triangle singularity from the dynamic pole structure, the simple pole configurations shall suffice. Our next step is to then generate an equal number of training datasets for each of these classifications.
@@
As previously discussed, there are multiple triangle diagrams that can produce enhancements within our energy region. For our task, we only consider the two loops, shown in Figure 3(a) and 3(b), that can produce a peak exactly at 4312 MeV so that our trained DNN can later match it to the experimental data. Inclusion of other relevant loops may be done to expand the training dataset but is not mandatory. We highlight once again that it is sufficient to use these possibly unrealistic diagrams, as we aim to test the likelihood of these mechanisms through the DNN. See section II for the detailed discussion. Moreover, we have shown that by varying the masses of the intermediate particles, one can construct enhancements at different locations. For this, we can consider the uncertainty in the masses of the hadrons involved, as well as attempt to vary the mass of particle 1, using equation (6), such that the produced enhancement still falls within our region of interest of [4200,4350] MeV. For triangle loop B, varying particle 1 means considering the different $\Lambda^{*}$ hyperon candidates as listed in PDG. For triangle loop A, it is simply adjusting the mass of the hypothetical $D_{s}^{**-}$ meson. It is also possible to vary the $\Lambda$ cut-off parameter and the $\epsilon$ decay width to control the strength and distribution of the resulting peaks. All the range of values used for these parameters are listed in Table 2. By randomly choosing a certain number of values for each of the parameters and taking all their combinations, we construct 5000 line shapes each for triangle loop A and triangle loop B. For convenience, we call these datasets T00A and T00B, respectively. Note that to mimic the experimental data, we include a projected phase space and an arbitrary background polynomial to the amplitudes following equation (8). A total of 10,000 training data is thus constructed for the triangle singularity.
@@
where energy $E_{\text{pole}}$ is in MeV. In addition to the main poles, we also produce distant background poles far from the scattering region, in order to simulate the smooth non-resonant background phase. Similarly, we include the projected phase space and arbitrary background polynomial to the amplitudes. By randomly generating 100 $\text{Re}\,E_{\text{pole}}$ and 100 $\text{Im}\,E_{\text{pole}}$ and taking their combinations, we construct 10,000 line shapes for each of the pole configurations listed in Table 3. We name these datasets as P01, P02, and P03. A sample and short discussion of all these training datasets can be found in appendix A. Finally, we compile all our data with their corresponding proper output labels and prepare for DNN training.
@@ IV.2 DNN architecture
The basic architecture of our DNN is shown in Figure 6. The input layer has 75 nodes for the energy points and 75 nodes for the intensities. In the output layer, we create 4 nodes such that each corresponds to a classification or label listed in Table 3. For the hidden layers, there are no strict guidelines as to how many layers or how many nodes should be constructed for optimal performance. Therefore, the task is to arbitrarily design different hidden layer architectures and compare their performances accordingly. In this study, we consider three different architectures enumerated in Table 4. The notation [XXX-…-XXX] denotes the number of layers and the number of nodes in each hidden layer. For example, [250-100] refers to an architecture with 250 nodes in the first hidden layer and 100 nodes in the second hidden layer. For a complete notation, we can also write 150-[250-100]-4 to include the input and output layers.
At the input layer, numerical values are entered through the nodes, and the main function of the DNN is to perform a linear combination of these inputs using certain weights and biases. This linear transformation is then fed to the nodes of the next layer. At each node of the succeeding layers, an activation function is equipped which shall assign a new numerical value to the node based on the received input, hence forming a new linear combination to be passed onto the next layer. For the hidden layer nodes, we use the rectified linear unit (ReLU) as the activation function, and the softmax for the output layer. As is standard, the softmax cross entropy is used as the cost function to be minimized Aggarwal (2018). This combination of linear transformations and non-linear effects of the activation functions allows the DNN to map the input space to the output space, which shall be optimized by adjusting the weights and biases during the training period. We follow the Chainer framework Tokui et al. (2015); Akiba et al. (2017); Tokui et al. (2019) in executing the construction and training of our DNN models. All the codes used in this study are accessible through our public repository in Co (2024).
Hidden layers [250-100], [250-100-50], [250-250-250]
Optimizers
Adam, AdaDelta, AdaGrad,
AMSGrad, AdaBound, AMSBound,
RMSprop, RMSpropGraves, SMORMS3
Table 4: List of hidden layer architectures and optimizers used in the study. For more information about the optimizers, see Cha .

### v3

@@ III Uniformization of $S$-matrix
Enhancements observed in scattering cross-sections most often correspond to poles in the scattering matrix. Different $S$-matrix parametrizations can be used to analyze a given experimental data. Perhaps the most popular one is the $K$-matrix parametrization since it automatically satisfies the unitarity of the $S$-matrix. However, identifying the location and Riemann sheet of the poles requires solving a polynomial in channel momenta. As such, their locations cannot be controlled to ensure a well-distributed representation of different possible pole locations. Moreover, it is usually the case that a pole can be placed on the desired Riemann sheet in the $K$-matrix approach, but is often accompanied by shadow poles which may appear in the physical Riemann sheet. Such pole configuration will result in a violation of causality and must not be included in our training dataset. It is, therefore, more practical to use a parametrization in which poles can be controlled directly. It turns out that this problem can be addressed by a more rigorous approach called uniformization Santos and Sombillo (2023).
@@
where $\epsilon=m_{1}+m_{2}$ is the threshold, and $\mu$ is the reduced mass. Notice that even though we write $p$ in terms of $\sqrt{s}$, there are still branch point singularities due to the existence of square root cross terms. To resolve this, we introduce the uniformized momentum $q_{i}$ to simplify the scattering energy, giving us
$s=q^{2}+\epsilon^{2},$ (11)
and
$s=q_{1}^{2}+\epsilon_{1}^{2}=q_{2}^{2}+\epsilon_{2}^{2},$ (12)
for a two-channel scattering.
The two-channel scattering can be described by a $2\times 2$ uniformized $S$-matrix with elements constructed using a Jost function $D$ and given by
$S_{11}(q_{1},q_{2})$ $=\dfrac{D(-q_{1},q_{2})}{D(q_{1},q_{2})},$ (13)
@@
which can be derived by utilizing the two expressions in equation (12). We can then construct a Jost function with the form
@@
The factor $(\omega-\omega_{\text{pole}})$ produces the poles in the $S$-matrix. The $1/\omega^{2}$ prohibits very energetic initial states from having no final states after the interaction as this is unphysical. The pole $\omega_{\text{reg}}$ ensures that the diagonal elements go to unity ($S_{ii}\rightarrow 1$) when the initial states have very large momentum ($\omega\rightarrow\infty$), which is a well-known property in potential scattering theory Taylor (1972). It is controlled such that it is far from the relevant pole and does not affect the interpretation of the enhancements. To satisfy hermiticity below the lowest threshold, there must be a pole in the negative complex conjugate of $\omega_{\text{pole}}$, that is why we have the factors ($\omega+\omega_{\text{pole}}^{*}$)($\omega+\omega_{\text{reg}}^{*}$).
Hence, by using uniformization, one has the freedom to adjust the position of the pole and its regulator to their desired Riemann sheets while preserving causality. In two-channel scattering, there are four Riemann sheets of the complex energy plane. The first Riemann sheet named [$tt$]-sheet is called the physical sheet where the scattering region is located. There should be no poles in the physical sheet to obey causality van Kampen (1953a, b). The other three Riemann sheets, namely [$bt$], [$bb$], and [$tb$]-sheets, are the unphysical sheets where poles located can be associated with the presence of an unstable quantum state in the reaction involved. One can say that the pole is located in the [$bt$]-sheet when the peak is below the second threshold, and a pole is located in the [$tb$]-sheet when the peak is located exactly at the threshold. These pole structures, and their manifestation on the physical region shown in Fig. 4, are useful in the description of the hadron’s internal structure as demonstrated in the pole-counting argument by Morgan and Pennington Morgan (1992); Morgan and Pennington (1991, 1993). The location of the Riemann sheets is shown in Table 1. The pole regulator in $D(\omega)$ has the form $\omega_{\text{reg}}=\exp(-i\pi/2)/|\omega_{\text{pole}}|$ where the phase factor forces the regulator to be in either the farther [$bb$]-sheet or the [$tb$]-sheet below the lowest threshold.
Figure 4: Expected line shapes arising from isolated poles in one of the unphysical Riemann sheets. The black vertical dotted line is the second threshold. The energy poles were generated with real part exactly at the second threshold and imaginary part of $-10$ MeV.
@@
The $J/\psi p$ scattering is a coupled-channel scattering problem where the first channel is set to always be $J/\psi p$ and the second channel is $\Sigma_{c}^{+}\bar{D}^{0}$ as this is the prominent threshold near the $P_{\psi}^{N}(4312)^{+}$ signal. One can construct its uniformized $S$-matrix where the elements $S_{11}$ describes the transition of $J/\psi p\rightarrow J/\psi p$, $S_{12}$ for $J/\psi p\rightarrow\Sigma_{c}^{+}\bar{D}^{0}$, $S_{21}$ for $\Sigma_{c}^{+}\bar{D}^{0}\rightarrow J/\psi p$, and $S_{22}$ for $\Sigma_{c}^{+}\bar{D}^{0}\rightarrow\Sigma_{c}^{+}\bar{D}^{0}$. A compact state arises when there are poles both in the [$bt$] and [$bb$]-sheets, corresponding to the mechanism in Fig. 3(c). By placing two identical energy poles on these sheets (i.e. the poles are equidistant from the relevant threshold), one can interpret the fitted enhancement as a pure compact state similar to fitting a Breit-Wigner amplitude Baru et al. (2004). On the other hand, a quasi-bound state interpretation is possible whenever there is an isolated pole in the $[bt]$-sheet with strong coupling to the higher channel, corresponding to Fig. 3(d). However, this interpretation should be examined carefully since a virtual state of the higher channel can move to the second Riemann sheet due to strong channel-coupling Fernández-Ramírez et al. (2019); Frazer and Hendry (1964); Pearce and Gibson (1989). It must be pointed out that bound and virtual states fall under the same general classification of hadronic molecules as demonstrated in Matuschek et al. (2021). In any case, any interpretation should be supported by the corresponding pole trajectory, which can only be done in a model-dependent way Badalyan et al. (1982); Baru et al. (2004); Hanhart et al. (2014).
@@
Through the uniformization scheme, one can easily construct pole-based amplitudes that correspond to the interpretations discussed. This is demonstrated in Fig. 5. We assign an isolated pole with energy (in units of MeV) at $4317.49-7j$ in the [$bt$]-sheet to construct an amplitude representing a possible hadronic molecule, while we assign one pole each at $4311.90-7j$ in both the [$bt$] and [$bb$]-sheets representing a compact pentaquark state. An amplitude is generally complex-valued with a real and imaginary part, but for our purposes, it is sufficient to consider its modulus square which corresponds to the intensity. Therefore, plotting the $|T_{11}|^{2}$ ($T$-matrix) contribution, where $S=1-2iT$, using the uniformized poles and following equation (8), we see that the two similar line shapes fit the LHCb experimental data, both with a prominent peak at 4312 MeV. If we compare our demonstrations in Fig. 5 and 2, we notice that the pole-based enhancements have a seemingly narrower and asymmetric width compared to that produced by a triangle singularity, and can fit the experimental error bars better especially near the $\Sigma_{c}^{+}\bar{D}^{0}$ threshold. Such asymmetry in the line shape is one of the characteristic features whenever the hadrons in the higher mass channel form a hadronic molecule. We take note of these observations as we move on to constructing the training dataset for our DNN.
@@ IV.1 Generation of training datasets
The experimental data is an invariant mass distribution containing a set of points specified by the energy value (horizontal axis) and the intensities (weighted candidates in the vertical axis). It is important to consider the energy values as part of our input features since they are expected to be smeared out (not equidistant) due to resolution effects of the particle detector. Thus, our goal is to construct a DNN model that accepts two-dimensional line shapes as input features and maps them to output targets corresponding to the possible classifications. Since we will eventually apply the trained DNN on interpreting the experimental data, the training dataset must follow the general characteristics of the experimental data (similar to our demonstrations in Fig. 2 and 5). In the LHCb data LHCb Collaboration (2019), the energy axis is divided into bins of 2 MeV each. For our training dataset, we can generate amplitudes on randomly spaced energy points, where one representative energy value is randomly chosen for each bin. We use a uniform distribution for this task. Moreover, we restrict our region of interest to the energy region of 4200 MeV to 4350 MeV, in order to avoid other prominent structures aside from the $P_{\psi}^{N}(4312)^{+}$. This region consists of 75 energy bins, implying that our line shapes shall have 75 points each for the energy value and weighted candidates. This totals to 150 input nodes, as presented in Fig. 6.
@@
For the output targets, we create labels for each of the possible classifications, as listed in Table 3. The triangle singularity takes up the first label, while the rest are possible pole configurations. As previously stated, 1 pole in [$bt$] may be associated to the hadronic molecule interpretation, 1 pole in [$tb$] corresponds to a virtual state, and 1 pole in [$bt$] and 1 pole in [$bb$] can accommodate a compact pentaquark interpretation. Refer to section III for the detailed discussion. It is possible to include more sophisticated pole configurations, such as considering 2 or more poles in each of the sheets as done in Sombillo et al. (2021a). However, for the purposes of our task which is mainly to distinguish triangle singularity from the dynamic pole structure, the simple pole configurations shall suffice. Our next step is to then generate an equal number of training datasets for each of these classifications.
@@
As previously discussed, there are multiple triangle diagrams that can produce enhancements within our energy region. For our task, we only consider the two loops, shown in Fig. 3(a) and 3(b), that can produce a peak exactly at 4312 MeV so that our trained DNN can later match it to the experimental data. Inclusion of other relevant loops may be done to expand the training dataset but is not mandatory. We highlight once again that it is sufficient to use these possibly unphysical diagrams, as we aim to test the likelihood of these mechanisms through the DNN. Moreover, we have shown that by varying the masses of the intermediate particles, one can construct enhancements at different locations. For this, we can consider the uncertainty in the masses of the hadrons involved, as well as vary the mass of particle 1, using equation (6), such that the produced enhancement still falls within our region of interest of [4200,4350] MeV. For triangle loop B, varying particle 1 means considering the different $\Lambda^{*}$ hyperon candidates as listed in PDG. For triangle loop A, it is simply adjusting the mass of the hypothetical $D_{s}^{**-}$ meson. It is also possible to vary the $\Lambda$ cut-off parameter and the $\epsilon$ decay width to control the strength and distribution of the resulting peaks.
Furthermore, it is important to recognize the possibility of producing cusp around the relevant threshold in the triangle singularity. In principle, one cannot separate the possible appearance of cusp in the triangle mechanism due to the presence of two-hadron interaction leading to the final state. In order to ensure that the triangle singularity is the dominant structure in the line shape, we ensure that the cusp effect is small by making the cut-off parameter very large. This means that the poles produced by the effective meson-baryon interaction is already far from the relevant scattering region. The cusp at the relevant threshold will still be present but it is negligible in comparison to the triangle enhancement.
All the range of values used for these parameters are listed in Table 2. By randomly choosing a certain number of values for each of the parameters and taking all their combinations, we construct 5000 line shapes each for triangle loop A and triangle loop B. For convenience, we call these datasets T00A and T00B, respectively. Note that to mimic the experimental data, we include a projected phase space and an arbitrary background polynomial (noise) to the amplitudes following equation (8). A total of 10,000 training data is thus constructed for the triangle singularity.
@@
where energy $E_{\text{pole}}$ is in MeV and $T_{2}$ is the $\Sigma_{c}^{+}\bar{D}^{0}$ threshold energy. In addition to the main poles, we also produce distant background poles far from the scattering region, in order to simulate the smooth non-resonant background phase. Similarly, we include the projected phase space and arbitrary background polynomial to the amplitudes. By randomly generating 100 $\text{Re}\,E_{\text{pole}}$ and 100 $\text{Im}\,E_{\text{pole}}$ and taking their combinations, we construct 10,000 line shapes for each of the pole configurations listed in Table 3. We name these datasets as P01, P02, and P03. A sample and short discussion of all these training datasets can be found in appendix A. Finally, we compile all our data with their corresponding proper output labels and prepare for DNN training.
@@ IV.2 DNN architecture
The basic architecture of our DNN is shown in Fig. 6. The input layer has 75 nodes for the energy points and 75 nodes for the intensities. In the output layer, we create 4 nodes such that each corresponds to a classification or label listed in Table 3. In this study, we consider three different architectures enumerated in Table 4. The notation [XXX-…-XXX] denotes the number of hidden layers and the number of nodes in each hidden layer. For example, [250-100] refers to an architecture with 250 nodes in the first hidden layer and 100 nodes in the second hidden layer. For a complete notation, we can also write 150-[250-100]-4 to include the input and output layers. For the hidden layers, we utilize the different DNN architectures used in Sombillo et al. (2021b) which are already tested and validated for the bound-virtual classification of single-channel scattering. For a more rigorous neural network architecture search, one may utilize the Bayesian optimization Zoph and Le (2017); Snoek et al. (2012); Frazier (2018). In the present study, the architectures shown in Table 4 should suffice for our purpose.
The basic operation of a typical feed-forward neural network is described as follows: A linear combination is formed using the node values of the preceding layer and the present weight and biases of the network. Except for the input layer, the hidden and the output layers are further equipped with activation functions which makes the mapping non-linear. We use the rectified linear unit (ReLU) as the activation function, and the softmax for the output layer. As is standard in a classification problem, the softmax cross entropy is used as the cost function to be minimized Aggarwal (2018). This combination of linear transformations and non-linear effects of the activation functions allows the DNN to map the input space to the output space, which are further improved by adjusting the weights and biases during the training iteration. We follow the Chainer framework Tokui et al. (2015); Akiba et al. (2017); Tokui et al. (2019) in executing the construction and training of our DNN models. All the codes used in this study are accessible through our public repository in Co (2024).
DNN model Optimizer and architecture
DNN 1 AdaGrad: 150-[250-100]-4
DNN 2 AdaGrad: 150-[250-100-50]-4
DNN 3 AdaGrad: 150-[250-250-250]-4
DNN 4 AMSGrad: 150-[250-100]-4
DNN 5 AMSGrad: 150-[250-100-50]-4
DNN 6 AMSGrad: 150-[250-250-250]-4
DNN 7 SMORMS3: 150-[250-100]-4
DNN 8 SMORMS3: 150-[250-100-50]-4
DNN 9 SMORMS3: 150-[250-250-250]-4
Table 4: DNN models used in the final analysis.